OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

Substitution means that you must solve for one variable in one equation ans then replace the value on the other. from the first equation we get y=1-2x Then replacing on the othe we get 4x+2(1-2x)=-2 That is a one variable equation.

OpenStudy (anonymous):

2x + y = 1 ---> y = 1 - 2x now sub 1 - 2x in for y in the other equation... 4x + 2y = - 2 4x + 2(1 - 2x) = - 2 (distribute the 2 through the parenthesis) 4x + 2 - 4x = - 2 4x - 4x = -2 -2 0 = - 4 This problem has no solution